OpenStudy (anonymous):

Fill in the missing value in the ordered pair for the equation 5x - 2y = 13. (3, _)

OpenStudy (anonymous):

First, you should rewrite the equation into y = f(x) and then substitute domain value to get range value :D

OpenStudy (anonymous):

5x-2y=13, since you already have the answer for the x coordinate (3) you can plug that into the equation: 5(3)-2y=13. Multiply 5 by 3. 15-2y=13 Then to find out the y value, move everything of the same value over to their own sides. -2y=13-15 Simplify. -2y=-2 Because they are both negatives and an x cannot be a negative, divide by -1 to make each positive. 2y=2 y=1 If you want to check the answer to make sure this is correct for the x also, do the same with the y value you just found. 5x-2y=13 5x-2(1)=13 5x-2=13 5x=13+2 5x=15 x=15/5 x=3. Therefore, (3,1) are the (x,y) values.